OpenStudy (anonymous):

find the following assume that a can represent any real number. ^4sqrt(8a)^4

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[4]{(8a)^4}=|8a| = |8||a| = 8|a|\] So \[\Large \sqrt[4]{(8a)^4} = 8|a|\]